Show that the lines

r·(4i+3j)−1=0, r·(−2i+j)−1=0, r·(i+2j)−1=0 ̃ ̃ ̃ ̃ ̃ ̃ ̃ ̃ ̃

are concurrent.


1
Expert's answer
2021-02-24T06:08:29-0500

Solution:

Comparing each of them with "\vec r.(a\hat i+b\hat j)+c=0"

we get,

"a_1=4, b_1=3,c_1=-1
\\ a_2=-2, b_2=1,c_2=-1
\\ a_3=1,b_3=2,c_3=-1"

For these lines to be concurrent, they must satisfy following:

"\begin{vmatrix}
a_1 & b_1 & c_1 \\ 
a_2 & b_2 & c_2 \\ 
a_3 & b_3 & c_3 \\ 
\end{vmatrix} =0"

Take LHS

"=\begin{vmatrix}
a_1 & b_1 & c_1 \\ 
a_2 & b_2 & c_2 \\ 
a_3 & b_3 & c_3 \\ 
\end{vmatrix}"

Putting above values

"=\begin{vmatrix}
4 & 3 & -1 \\ 
-2 & 1 & -1 \\ 
1 & 2 & -1 \\ 
\end{vmatrix}"

"=4\begin{vmatrix}
1 & -1 \\ 
2 & -1 \\ 
\end{vmatrix}-3\begin{vmatrix}
-2 & -1 \\ 
1 & -1 \\ 
\end{vmatrix}+(-1)\begin{vmatrix}
-2 & 1 \\ 
1 & 2 \\ 
\end{vmatrix}"

"=4(-1+2)-3(2+1)-1(-4-1)
\\ =4-9+5
\\ =9-9
\\ =0"

"=" RHS

Thus, above lines are concurrent.